OpenStudy (anonymous):

PLEASE HELP ASAP!!! On a movie set, an archway is modeled by the equation y = -0.5x2 + 3x, where y is the height in feet and x is the horizontal distance in feet. A laser is directed at the archway at an angle modeled by the equation -0.5x + 2.42y = 7.65 such that the beam crosses the archway at points A and B. At what height from the ground are the points A and B?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

solve for y from -0.5x + 2.42y = 7.65 then y = y

OpenStudy (anonymous):

I dont understand

jigglypuff314 (jigglypuff314):

solve for y from -0.5x + 2.42y = 7.65 what do you get?

OpenStudy (anonymous):

1y=0.21x+3.16

jigglypuff314 (jigglypuff314):

yes then since y = y -0.5x^2 + 3x = 0.21x+3.16

OpenStudy (anonymous):

the choices are: a:1.5ft, 3.5 ft b:1.5 ft, 4 ft. c: 3.5ft, 4 ft d: 4 ft, 4 ft

OpenStudy (anonymous):

Do you know what the answer is?

jigglypuff314 (jigglypuff314):

-0.5x^2 + 3x = 0.21x+3.16 -0.21x -0.21x -0.5x^2 + 2.79x - 3.16 = 0 solve for x... I suppose you can use the quadratic thing or wolframalpha.com

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

What do i do next?

jigglypuff314 (jigglypuff314):

there are supposed to be two answers

OpenStudy (anonymous):

yeah. the x and the y

OpenStudy (anonymous):

I posted the answer choices. I dont know which one to choose

OpenStudy (anonymous):

OH. THANKS LOL